# Multi-wavelength Emission from Jets and Magnetically Arrested Disks in Nearby Radio Galaxies: Application to M87

### v1

@@ 5.3.1 Neutrino luminosity from Jet
We estimate neutrino flux produced by the photomeson production. First, we estimate the maximum Lorentz factor of the protons. At the initial stage of the plasma entrainment, protons are a subdominant component of the internal energy (i.e., $n_{e}m_{e}\gtrsim n_{p}m_{p}$). Then, the proton magnetization parameter is $\sigma_{p}\simeq\sigma_{\pm}$, and the magnetic reconnection can achieve $\gamma_{p}\sim(\delta B/B)^{2}\xi\sigma_{\pm}\simeq 10^{7}$, where $\gamma_{p}$ is the proton Lorentz factor. Neutrino energy is estimated to be $E_{\nu}\approx 0.05E_{p}$. Then, $E_{\nu}\simeq 5.0\times 10^{14}~{}\rm eV$ for $E_{p}\simeq 1.0\times 10^{16}~{}\rm eV$. For the photomeson production, the protons interact with the photons of $E_{\gamma}\simeq 0.3~{}\rm GeV$ in the proton rest frame for the delta resonance approximation. Then, the protons of $\gamma_{p}\simeq 10^{7}$ interact with $E_{\gamma}\simeq 30~{}\rm eV$, which is in the UV band. We estimate the number density of the UV photons inside the jet as $n_{\rm UV}\approx F_{\rm UV}D^{2}/(R_{j}^{2}c\epsilon_{\rm UV})\simeq 2.8\times 10^{5}\ \rm cm^{-3}$, where $F_{\rm UV}=1.3\times 10^{-12}\ ~{}\rm erg~{}s^{-1}~{}cm^{-2}$ is the flux of the UV photons, and $\epsilon_{\rm UV}\simeq 30\ \rm eV$ is the UV photon energy. We estimate the photomeson production timescale as $t_{p\gamma}=1/(n_{\rm UV}\sigma_{p\gamma}\kappa_{p\gamma}c)\simeq 1.2\times 10^{12}~{}\rm s$, where $\sigma_{p\gamma}\approx 0.5~{}\rm mb$ is the cross section and $\kappa_{p\gamma}\approx 0.2$ is the inelasticity of the photomeson production, respectively. By equation (12), we estimate the proton luminosity as $L_{p}=L_{j}/(1+0.5\sigma_{\rm dis})-L_{e}\simeq L_{j}$. Then, we optimistically estimate the luminosity of the neutrino as $L_{\nu}\approx f_{p\gamma}L_{p}\simeq 5.8\times 10^{37}~{}\rm erg~{}s^{-1}$, where $f_{p\gamma}=t_{\rm loss}/t_{p\gamma}\simeq 1.2\times 10^{-7}$ is the pion production efficiency. Regarding the loss timescale, diffusive escape is dominant, compared to the adiabatic cooling, in the relevant energy (see bottom panel of Figure 4), and the proton synchrotron cooling timescale is negligibly long. The neutrino flux is estimated to be $F_{\nu}=L_{\nu}/(4\pi D^{2})\simeq 1.7\times 10^{-15}\ \rm erg~{}s^{-1}~{}cm^{-2}$, which is challenging to detect with near-future neutrino detectors.
@@ 5.3.2 Neutrino luminosity from MAD
Next, we estimate the neutrino emission from MADs. As shown in Figure 3, the proton maximum Lorentz factor is estimated to be $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$. We estimate the proton number density to be $n_{p}\approx\rho/m_{p}\simeq 4.1\times 10^{4}\ \rm cm^{-3}$, and then, the $pp$ inelastic collision timescale is estimated to be $t_{pp}\simeq 2.7\times 10^{10}\ \rm s$. The $pp$ inelastic collision timescale is longer than the diffusion timescale $t_{\rm diff}\simeq 2.2\times 10^{6}\ \rm s$. $t_{pp}$ is also much longer than photomeson production timescale, $t_{p\gamma}$, as shown below. Thus, protons emit neutrinos via photomeson production.
We estimate the neutrino flux of the photomeson production in the same way as the emission from the jet. The protons of $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$ interact with $6.3\times 10^{-2}$ eV photons, and we estimate the number density of the infrared photons as $n_{\rm IR}\approx F_{\rm IR}D^{2}/(R_{d}^{2}c\epsilon_{\rm IR})\simeq 4.9\times 10^{8}\ \rm cm^{-3}$, where $F_{\rm IR}=4.7\times 10^{-14}~{}\rm erg~{}s^{-1}~{}cm^{-2}$ is the flux of the infrared photons emitted by the Comptonization process and $\epsilon_{\rm IR}\simeq 6.3\times 10^{-2}\ \rm eV$ is the infrared photon energy. Then, the photomeson production timescale is estimated to be $t_{p\gamma}\simeq 6.9\times 10^{8}~{}\rm s$. The nonthermal proton luminosity of the MAD is estimated to be $L_{p}(\gamma_{p,\rm max})\approx\epsilon_{\rm NT}\epsilon_{\rm dis}\dot{M}c^{2}\exp(-1)\simeq 8.7\times 10^{41}\ \rm erg~{}s^{-1}$, where $\epsilon_{\rm NT}=0.33$ is the fraction of the nonthermal particle energy production rate to the dissipation rate and $\epsilon_{\rm dis}=0.15$ is the fraction of the dissipation to the accretion rate. We estimate the luminosity of the neutrino and the pion production efficiency as $L_{\nu}\approx f_{p\gamma}L_{p}(\gamma_{p,\rm max})\simeq 1.3\times 10^{39}~{}\rm erg~{}s^{-1}$ and $f_{p\gamma}\simeq 1.5\times 10^{-3}$, respectively. We estimate the neutrino flux from the MAD to be $F_{\nu}\simeq 3.7\times 10^{-14}\ \rm erg~{}s^{-1}~{}cm^{-2}$. We also numerically compute the neutrino flux using the method given in Section 3 of Kimura & Toma (2020), which is consistent with our analytic estimate within a factor of 1.5. Protons of $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$ effectively produce the neutrinos of $2.4\times 10^{17}~{}\rm eV$. To observe the sub-EeV neutrinos, experiments that detect radio signals from the neutrino-induced particle shower, such as GRAND (Álvarez-Muñiz et al., 2019), RNO (Aguilar et al., 2021), and Gen2-Radio (Hallmann et al., 2021), are important.
@@ 6 Conclusion
We construct the two-zone multi-wavelength emission model, the ‘Jet-MAD model’, that takes into account the plasma loading to the jet based on the model proposed by Kimura et al. (2022) and Chen et al. (2023). In the MAD, the MHD turbulence heats up the thermal electrons that emit thermal synchrotron radiation. The MHD turbulence also accelerates the nonthermal protons, which emit multi-wavelength photons via synchrotron radiation. For the jet, we consider that the velocity difference between the jet and the ambient gas generates turbulence via the KHI. This turbulence induces the entrainment of the ambient gas into the jet. The turbulence also triggers the magnetic reconnection inside the jet, accelerating the electrons that efficiently emit multi-wavelength photons. In this scenario, the Jet-MAD model can explain the simultaneous multi-wavelength data for M87 with a parameter set given in Table 1. Synchrotron radiation from the jet explains the optical, UV, and X-ray data, and synchrotron radiation of the protons from the MAD explains the GeV gamma-ray data.
@@
These results suggest that the jet consists of the electron-positron pairs produced by the Breit-Wheeler process at the jet base and the electron-proton plasma entrained from the ambient gas via the KHI. These results also suggest that the plasma entrainment is strong enough to satisfy $(\delta B/B)^{2}\sigma_{\rm ent}<1$ and the magnetic reconnection scenario likely works as the nonthermal particle acceleration mechanism inside the jet. Our conclusion remains the same even if we have a power-law tail component for $\gamma>\gamma_{\rm cut}$.
@@
Time variability helps us to test our model. Our model predicts that optical and X-rays are emitted from the jet, while submillimeter and gamma rays are emitted from the MAD. Based on the size of each emission region, we roughly estimate the variability timescales to be $10^{6}\;$s for the jet and $10^{5}\;$s for the MAD. Thus, pointing multi-wavelength instruments to M87 for more than one day is needed to test our model. Then, high-time resolution observations at least higher than $10^{5}\;$s such as submillimeter observation with ALMA, infrared and optical observations with JWST and HST, X-ray observations with CHANDRA or Swift XRT, and TeV gamma-ray observations with MAGIC, HESS, and CTA (Actis et al., 2011), are desired. However, the sensitivity of the Fermi-LAT which observes GeV gamma rays is insufficient to detect the time variability (Atwood et al., 2009). Our model predicts that submillimeter and TeV gamma-ray light curves strongly correlate with day-scale time variability. Also, optical and X-ray light curves should correlate strongly with week-scale variability. Future multi-wavelength observation campaigns will provide a test of our model.
@@ acknowledgments
We would like to thank Wataru Ishizaki and Hayk Hakobyan for useful comments. Discussions for this paper were performed many times at Science Lounge of FRIS CoRE. This work is partly supported by JSPS KAKENHI grant Nos. 22K14028, 21H04487, and 23H04899 (S. S. K.). This work is also supported by JST SPRING, grant No. JPMJFS2114 (R.K.). S.S.K. acknowledges support by the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXTs Strategic Professional Development Program for Young Researchers.
Appendix A Magnetization parameter for power-law tail distribution

### v2

@@ 5.3.1 Neutrino luminosity from Jet
We estimate neutrino flux produced by the photomeson production. First, we estimate the maximum Lorentz factor of the protons. At the initial stage of the plasma entrainment, protons are a subdominant component of the internal energy (i.e., $n_{e}m_{e}\gtrsim n_{p}m_{p}$). Then, the proton magnetization parameter is $\sigma_{p}\simeq\sigma_{\pm}$, and the magnetic reconnection can achieve $\gamma_{p}\sim(\delta B/B)^{2}\xi\sigma_{\pm}\simeq 10^{7}$, where $\gamma_{p}$ is the proton Lorentz factor. Neutrino energy is estimated to be $E_{\nu}\approx 0.05E_{p}$. Then, $E_{\nu}\simeq 5.0\times 10^{14}~{}\rm eV$ for $E_{p}\simeq 1.0\times 10^{16}~{}\rm eV$. For the photomeson production, the protons interact with the photons of $E_{\gamma}\simeq 0.3~{}\rm GeV$ in the proton rest frame for the delta resonance approximation. Then, the protons of $\gamma_{p}\simeq 10^{7}$ interact with $E_{\gamma}\simeq 30~{}\rm eV$, which is in the UV band. We estimate the number density of the UV photons inside the jet as $n_{\rm UV}\approx F_{\rm UV}D^{2}/(R_{j}^{2}c\epsilon_{\rm UV})\simeq 2.8\times 10^{5}\ \rm cm^{-3}$, where $F_{\rm UV}=1.3\times 10^{-12}\ ~{}\rm erg~{}s^{-1}~{}cm^{-2}$ is the flux of the UV photons, and $\epsilon_{\rm UV}\simeq 30\ \rm eV$ is the UV photon energy. We estimate the photomeson production timescale as $t_{p\gamma}=1/(n_{\rm UV}\sigma_{p\gamma}\kappa_{p\gamma}c)\simeq 1.2\times 10^{12}~{}\rm s$, where $\sigma_{p\gamma}\approx 0.5~{}\rm mb$ is the cross section and $\kappa_{p\gamma}\approx 0.2$ is the inelasticity of the photomeson production, respectively. By equation (6), we estimate the proton luminosity as $L_{p}=L_{j}/(1+0.5\sigma_{\rm dis})-L_{e}\simeq L_{j}$. Then, we optimistically estimate the luminosity of the neutrino as $L_{\nu}\approx f_{p\gamma}L_{p}\simeq 5.8\times 10^{37}~{}\rm erg~{}s^{-1}$, where $f_{p\gamma}=t_{\rm loss}/t_{p\gamma}\simeq 1.2\times 10^{-7}$ is the pion production efficiency. Regarding the loss timescale, diffusive escape is dominant, compared to the adiabatic cooling, in the relevant energy (see Figure 4), and the proton synchrotron cooling timescale is negligibly long. The neutrino flux is estimated to be $F_{\nu}=L_{\nu}/(4\pi D^{2})\simeq 1.7\times 10^{-15}\ \rm erg~{}s^{-1}~{}cm^{-2}$, which is challenging to detect with near-future neutrino detectors.
@@ 5.3.2 Neutrino luminosity from MAD
Next, we estimate the neutrino emission from MADs. As shown in Figure 2, the proton maximum Lorentz factor is estimated to be $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$. We estimate the proton number density to be $n_{p}\approx\rho/m_{p}\simeq 4.1\times 10^{4}\ \rm cm^{-3}$, where $\rho$ is the mass density, and then, the $pp$ inelastic collision timescale is estimated to be $t_{pp}\simeq 2.7\times 10^{10}\ \rm s$. The $pp$ inelastic collision timescale is longer than the diffusion timescale $t_{\rm diff}\simeq 2.2\times 10^{6}\ \rm s$. $t_{pp}$ is also much longer than photomeson production timescale, $t_{p\gamma}$, as shown below. Thus, protons emit neutrinos via photomeson production.
We estimate the neutrino flux of the photomeson production in the same way as the emission from the jet. The protons of $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$ interact with $6.3\times 10^{-2}$ eV photons, and we estimate the number density of the infrared photons as $n_{\rm IR}\approx F_{\rm IR}D^{2}/(R_{d}^{2}c\epsilon_{\rm IR})\simeq 4.9\times 10^{8}\ \rm cm^{-3}$, where $F_{\rm IR}=4.7\times 10^{-14}~{}\rm erg~{}s^{-1}~{}cm^{-2}$ is the flux of the infrared photons emitted by the Comptonization process and $\epsilon_{\rm IR}\simeq 6.3\times 10^{-2}\ \rm eV$ is the infrared photon energy. Then, the photomeson production timescale is estimated to be $t_{p\gamma}\simeq 6.9\times 10^{8}~{}\rm s$. The nonthermal proton luminosity of the MAD is estimated to be $L_{p}(\gamma_{p,\rm max})\approx\epsilon_{\rm NT}\epsilon_{\rm dis}\dot{M}c^{2}\exp(-1)\simeq 8.7\times 10^{41}\ \rm erg~{}s^{-1}$ (see Table 1). We estimate the luminosity of the neutrino and the pion production efficiency as $L_{\nu}\approx f_{p\gamma}L_{p}(\gamma_{p,\rm max})\simeq 1.3\times 10^{39}~{}\rm erg~{}s^{-1}$ and $f_{p\gamma}\simeq 1.5\times 10^{-3}$, respectively. We estimate the neutrino flux from the MAD to be $F_{\nu}\simeq 3.7\times 10^{-14}\ \rm erg~{}s^{-1}~{}cm^{-2}$. We also numerically compute the neutrino flux using the method given in Section 3 of Kimura & Toma (2020), which is consistent with our analytic estimate within a factor of 1.5. Protons of $\gamma_{p,\rm max}\simeq 4.8\times 10^{9}$ effectively produce the neutrinos of $2.4\times 10^{17}~{}\rm eV$. To observe the sub-EeV neutrinos, experiments that detect radio signals from the neutrino-induced particle shower, such as GRAND (Álvarez-Muñiz et al., 2019), RNO (Aguilar et al., 2021), and Gen2-Radio (Hallmann et al., 2021), are important.
@@ 5.4 Effect of the pitch angle anisotropy
5.5 Time variability analysis
Time variability helps us to test our model. Our model predicts that optical and X-rays are emitted from the jet, while mm/sub-mm and gamma rays are emitted from the MAD. Based on the size of each emission region, we roughly estimate the variability timescales to be $10^{6}\;$s for the jet and $10^{5}\;$s for the MAD. Thus, pointing multi-wavelength instruments to M87 for more than one day is needed to test our model. Then, high-time resolution observations at least higher than $10^{5}\;$s such as mm/sub-mm observation with ALMA, infrared and optical observations with JWST and HST, X-ray observations with CHANDRA or Swift XRT, and TeV gamma-ray observations with MAGIC, HESS, and CTA (Actis et al., 2011), are desired. However, the sensitivity of the Fermi-LAT, which observes the GeV gamma rays, is insufficient to detect the time variability (Atwood et al., 2009). Our model predicts that mm/sub-mm and TeV gamma-ray light curves strongly correlate with day-scale time variability. Also, optical and X-ray light curves should correlate strongly with week-scale variability. We should note that the time variability cannot distinguish between strong and weak plasma entrainments since we consider the same dissipation region inside the jet. Future multi-wavelength observation campaigns will provide a test of our model.
@@ 6 Conclusion
We construct the two-zone multi-wavelength emission model, the ‘Jet-MAD model’, that takes into account the plasma loading to the jet based on the model proposed by Kimura et al. (2022) and Chen et al. (2023). In the MAD, the MHD turbulence heats up the thermal electrons that emit thermal synchrotron radiation. The MHD turbulence also accelerates the nonthermal protons, which emit multi-wavelength photons via synchrotron radiation. For the jet, we consider that the velocity difference between the jet and the ambient gas generates turbulence via the KHI. This turbulence induces the entrainment of the ambient gas into the jet. The turbulence also triggers the magnetic reconnection inside the jet, accelerating the electrons that efficiently emit multi-wavelength photons. In this scenario, the Jet-MAD model can explain the simultaneous multi-wavelength data for M87 with a parameter set given in Table 1. Synchrotron radiation from the jet explains the optical, UV, and X-ray data, and synchrotron radiation of the thermal electrons and protons from the MAD explains the mm/sub-mm and GeV gamma-ray data, respectively.
@@
These results suggest that the jet consists of the electron-positron pairs produced by the Breit-Wheeler process at the jet base and the electron-proton plasma entrained from the ambient gas via the KHI. These results also suggest that the plasma entrainment is strong enough to satisfy $(\delta B/B)^{2}\sigma_{\rm ent}<1$ and the magnetic reconnection scenario likely works as the nonthermal particle acceleration mechanism inside the jet. Our conclusion remains the same even if we have a power-law tail component for $\gamma>\gamma_{\rm cut}$. Time variability analysis should test our model. Our model predicts a strong correlation between optical and X-ray variabilities, and the strong correlation should also appear in mm/sub-mm and gamma-ray variabilities.
@@
In this paper, we focus on the case study of M87. We can test our model by applying it to other radio galaxies, such as NGC 1275. This kind of test will make the applicability of our model more robust, but the detailed analysis and discussion will be left as future work.
@@ acknowledgments
We would like to thank Wataru Ishizaki and Hayk Hakobyan for their useful comments. The codes used for this study are available from R.K. or S.S.K. upon reasonable request. Discussions for this paper were performed many times at Science Lounge of FRIS CoRE. This work is partly supported by JSPS KAKENHI grant Nos. 22K14028, 21H04487, and 23H04899 (S. S. K.). This work is also supported by JST SPRING, grant No. JPMJFS2114 (R.K.). S.S.K. acknowledges support by the Tohoku Initiative for Fostering Global Researchers for Interdisciplinary Sciences (TI-FRIS) of MEXTs Strategic Professional Development Program for Young Researchers.
Appendix A Detailed calculation method in the MAD model
In this appendix, we explain the detailed calculation method in the MAD model. First, we explain how to determine the electron temperature in the MAD model. As explained in Section 2.1, we determine the temperature of the thermal electrons by equating the heating rate with the cooling rate:
$L_{\rm thermal}\approx\frac{Q_{i}}{Q_{e}}(1-\epsilon_{\rm NT})\epsilon_{\rm dis}\dot{M}c^{2},$ (A1)
where $L_{\rm thermal}$ is the luminosity of the photons emitted by the thermal electrons. As the electron heating rate, we use the description of Kawazura et al. (2020) which shows that when the compressive wave is dominant, the ratio of the electron-proton heating rate is approximated to be compressive-to-Alfvénic wave power ratio:
$\frac{Q_{i}}{Q_{e}}\approx\frac{P_{\rm compr}}{P_{\rm AW}},$ (A2)
where $P_{\rm compr}$ is the compressive wave power and $P_{\rm AW}$ is the Alfvénic wave power. The wave power ratio in the MAD has not been established yet. We give $Q_{i}/Q_{e}$ as a parameter and set it as 10. Since the accretion flow is in the MAD state, the plasma beta is likely lower than 1. Hence, $Q_{i}/Q_{e}\simeq 10$ corresponds to $R_{\rm low}\simeq 10$ in the $R_{\rm low}-R_{\rm high}$ prescription (see Mościbrodzka et al., 2016), and $R_{\rm low}\simeq 10$ is favored by the polarization observation of M87 (Event Horizon Telescope Collaboration et al., 2021b).
Next, we explain how to calculate the energy distributions of the nonthermal particles. To obtain the particle energy distributions, we solve the energy transport equation with the one-zone and steady-state approximations:
$-\frac{d}{dE_{i}}\left(\frac{E_{i}N_{E_{i}}}{t_{i,\rm cool}}\right)=\dot{N}_{E_{i},\rm inj}-\frac{N_{E_{i}}}{t_{\rm esc}},$ (A3)
where $N_{E_{i}}$ is the differential number spectrum, $t_{i,\rm cool}$ is the cooling time, $t_{\rm esc}$ is the escape time, $\dot{N}_{E_{i},\rm inj}$ is the injection terms, and $i$ is the particle species. We write the injection terms as
$\dot{N}_{E_{i},\rm inj}\approx\dot{N}_{0,\rm mad}\left(\frac{E_{i}}{E_{i,\rm cut}}\right)^{-s_{\rm inj}}\exp\left(-\frac{E_{i}}{E_{i,\rm cut}}\right).$ (A4)
[…]
$D_{pp}\approx\frac{p^{2}}{3\eta_{\rm turb}}\left(\frac{c}{H}\right)\left(\frac{V_{A}}{c}\right)^{2},$ (A5)
where $V_{A}\approx B_{d}/\sqrt{4\pi\rho}$ is the Alfvén velocity and $B_{d}$ is the magnetic field in the MAD. We estimate $B_{d}$ as
$B_{d}\approx\sqrt{\frac{8\pi\rho C_{s}^{2}}{\beta}}\simeq 62~{}\dot{m}_{-4}^{1/2}M_{9}^{-1/2}r_{1}^{-5/4}\alpha_{-0.5}^{-1/2}\beta_{-1}^{-1/2}~{}\rm G,$ (A6)
where $\rho\approx\dot{M}/(4\pi R_{d}HV_{R})$ is the mass density, $H\sim(C_{s}/V_{K})R_{d}$ is the scale height of the MADs, $C_{s}\approx V_{K}/2$ is the sound speed, $V_{R}\approx\alpha V_{K}$ is the radial velocity, and $V_{K}=\sqrt{GM_{\rm BH}/R_{d}}$ is the Kepler velocity (see also Kimura & Toma, 2020; Kuze et al., 2022). Then, we can write the acceleration timescale as
$t_{\rm acc}=\frac{p^{2}}{D_{pp}}\approx 3\eta_{\rm turb}\frac{H}{c}\left(\frac{c}{V_{A}}\right)^{2}.$ (A7)
We consider the relevant cooling processes (synchrotron, Bethe-Heitler, photomeson production), diffusive escape, and infall to the BH for the loss processes. We estimate the cooling timescales based on Kimura et al. (2019a). The synchrotron cooling timescale is estimated to be
$t_{p,\rm syn}=\frac{6\pi m_{e}c}{\gamma_{p}\sigma_{T}B_{d}^{2}}\left(\frac{m_{p}}{m_{e}}\right)^{3},$ (A8)
where $\sigma_{T}$ is the Thomson cross-section.
We estimate the photomeson production cooling timescale as
[…]
where $\bar{\epsilon}_{\rm th}\simeq 145~{}\rm MeV$ is the threshold energy for photomeson production and $\bar{\epsilon}_{\gamma}$ is the photon energy in proton rest frame. We estimate the cooling timescale of the Bethe-Heitler process, $t_{\rm BH}$, by Equation (A9) using $\sigma_{\rm BH}$ and $\kappa_{\rm BH}$ instead of $\sigma_{p\gamma}$ and $\kappa_{p\gamma}$, respectively. For $\sigma_{p\gamma}$ and $\kappa_{p\gamma}$, we use the fitting formulas given in Murase & Nagataki (2006), and for $\sigma_{\rm BH}$ and $\kappa_{\rm BH}$, we use those from Stepney & Guilbert (1983) and Chodorowski et al. (1992), respectively. We estimate the cooling timescale as $t_{\rm cool}^{-1}=t_{p,\rm syn}^{-1}+t_{p\gamma}^{-1}+t_{\rm BH}^{-1}$.
For diffusive escape, we estimate the timescale as $t_{\rm diff}=R_{d}^{2}/D_{zz}$, where $D_{zz}=(1/3)\eta_{\rm turb}cH$ is the diffusion coefficient in space. We estimate the infall timescale as $t_{\rm fall}=R_{d}/V_{R}$, and then, the total escape timescale is given as $t_{\rm esc}^{-1}=t_{\rm fall}^{-1}+t_{\rm diff}^{-1}$.
Appendix B Steady-state magnetization parameter in the reconnection region
In this appendix, we show the detailed calculation model for the steady-state magnetization parameter in the reconnection region. We consider that the magnetic reconnection in the BH magnetosphere accelerates the particles emitting the high-energy photons, and the emitted photons interact with each other by the Breit-Wheeler process, resulting in pair production. The magnetic reconnection will produce the power-law electron energy distribution with the index $p_{\rm inj}\sim 1$ (e.g., Sironi & Spitkovsky, 2014; Guo et al., 2014, 2015, 2016; Werner et al., 2016). These injected electrons rapidly cool as the cooling timescale is much shorter than the dynamical timescale (Kimura et al., 2022), resulting in the index of nonthermal electrons, $p\sim 2$. Then, the synchrotron spectrum is $dN/dE_{\gamma}\propto E_{\gamma}^{-1.5}$ for $E_{\gamma}<E_{\gamma,\rm cut,ms}$, where $E_{\gamma,\rm cut,ms}=\gamma_{\rm cut,ms}^{2}heB_{\rm rec}/(2\pi m_{e}c)$ is the photon energy emitted by the electrons with $\gamma_{\rm cut,ms}$ and $\gamma_{\rm cut,ms}$ is the cutoff Lorentz factor achieved by the magnetic reconnection in the BH magnetosphere. We assume the cutoff Lorentz factor as
$\gamma_{\rm cut,ms}=4\sigma_{\pm}.$ (B1)
The synchrotron photon spectrum has an exponential cutoff for $E_{\gamma}>E_{\gamma,\rm cut,ms}$.
$n_{\pm}=\dot{n}_{\pm}\frac{l_{\rm rec}}{c},$ (B2)
where $\dot{n}_{\pm}$ is the pair production rate. $\dot{n}_{\pm}$ depends on the photon spectrum, which in turn depends on the electron spectrum. Thus, $\dot{n}_{\pm}$ depends on $\sigma_{\pm}$ through Equation (B1). As shown in Equation (1), we estimate the steady-state magnetization parameter following the method of Chen et al. (2023).
Figure 9: Magnetization parameters for BH mass of $10^{9}M_{\odot}$ as a function of $\dot{m}$, calculated in the methods of Chen et al. (2023) (red solid line) and Kimura et al. (2022) (blue dashed line) and estimated by Equation (B3) (green dash-dotted line).
We show $\sigma_{\pm}$ for the BH mass of $10^{9}M_{\odot}$ as a function of $\dot{m}$ in Figure 9. As shown in Figure 9, $\sigma_{\pm}$ approximately depends on $\dot{m}^{-1/4}$. This comes from self-regulation. For $E_{\gamma,\rm cut,ms}\gg m_{e}c^{2}$, copious pairs are produced, and then, pair loading reduces $\sigma_{\pm}$. On the other hand, for $E_{\gamma,\rm cut,ms}\ll m_{e}c^{2}$, plasma loading stops, and the upstream plasma will escape in the dynamical time, which will increase $\sigma_{\pm}$. Thus, the peak of the synchrotron spectrum is adjusted to $E_{\gamma,\rm cut,ms}\approx 0.5m_{e}c^{2}$. This can be rewritten as
$\sigma_{\pm}\simeq 5.9\times 10^{4}M_{9}^{1/4}\dot{m}_{-4}^{-1/4}\Phi_{\rm rec,1.2}^{-1/2}.$ (B3)
Kimura et al. (2022) also estimate the magnetization parameter after pair loading (see their Equation (11)), but our estimate of $\sigma_{\pm}$ differs from that given in Kimura et al. (2022). They assume photon spectrum with an abrupt cutoff (i.e., no photons above $E_{\gamma,\rm cut,ms}$), which leads to different values of $\sigma_{\pm}$ and conclusions for such a range of mass accretion rates that we consider in this paper. There is another point that differs between ours and that given in Kimura et al. (2022). Ours takes into account the feedback by the pair loading when $m_{e}c^{2}<E_{\gamma,\rm cut,ms}<E_{\gamma,\rm rad}$, where $E_{\gamma,\rm rad}$ is the synchrotron photon energy for the burn-off limit (see Equation (5) in Kimura et al. (2022)), which help achieving self-regulation given in Equation (B3). In contrast, for the cases with a mass accretion rate as low as that in Sgr A*, the estimate given in Kimura et al. (2022) should be valid, as neither the feedback nor the tail distribution affects the pair loading process. Our model of pair loading assumes that $E_{\gamma,\rm cut,ms}$ determines the cutoff energy because of the sufficient pair loading, which contradicts the situation of low accretion rate cases where $E_{\gamma,\rm rad}<E_{\gamma,\rm cut,ms}$ is satisfied.
Appendix C Jet Lorenz factor and the size of the dissipation region
We estimate the jet Lorentz factor and Doppler factor as follows. The very long baseline interferometry (VLBI) observations for M87 estimate the jet width as $R\approx z^{2/3}$, where $z$ is the distance from the BH (Hada et al., 2013; Nakamura et al., 2018). Thus, we assume the size of the dissipation region as $R_{j}=10^{2}R_{g}$ to have the distance of the dissipation region $z_{j}\simeq 10^{3}R_{g}$. We set the jet Lorentz factor $\Gamma_{j}$ at the dissipation region as $\Gamma_{j}\beta_{j}\simeq 0.8$ based on the observation of the superluminal motion of the M87 jet with KaVA and VLBA (Park et al., 2019). The M87 jet inclination angle is observed as $\theta_{j}\approx 17\ \rm deg$ by the observation of jet and counter jet proper motions (Walker et al., 2018). Hence, we estimate the Doppler factor to be $\delta_{D}\approx 1.9$. The variability timescale is then estimated to be $t_{v}=R_{j}/(c\delta_{D})\simeq 1.6\times 10^{6}\;R_{j,2}\;{\rm s}$, which is consistent with the observed X-ray daily variability (Imazawa et al., 2021; EHT MWL Science Working Group et al., 2021). We calculate the photon spectra observed on Earth by $\nu_{\rm obs}L_{\nu,\rm obs}=\delta_{D}^{4}\nu_{\rm src}L_{\nu,\rm src}$, where $\nu_{\rm src}L_{\nu,\rm src}$ is the intrinsic photon spectrum obtained by the jet model.
Appendix D Cooling and escape timescales inside the jet
In this appendix, we explain how to estimate the cooling and escape timescales in the jet. As explained in Section 2.4, we consider the electrons cool through synchrotron, synchrotron self-Compton, and adiabatic expansion. We estimate the cooling timescale as $t_{\rm cool}^{-1}=t_{\rm syn}^{-1}+t_{\rm SSC}^{-1}+t_{\rm ad}^{-1}$.
We estimate the synchrotron cooling timescale to be $t_{\rm syn}=6\pi m_{e}c/(\gamma\sigma_{T}B_{\rm dis}^{2})$. We calculate the synchrotron self-Compton scattering cooling timescale in the way of subsection 2.6 of Blumenthal & Gould (1970). In this paper, we ignore the effect of the external photon field. The accretion rate is low for M87, and thus, the energy density of the external photon, such as thermal emission from a dust torus, is lower than that of the internal photon emitted by the energetic electrons.
We consider that the electrons are scattered by the KHI-induced turbulence. KHI injects the turbulence with the scale of the dissipation region, $R_{j}$, and cascades into the small scale. We assume that the spectrum of the turbulence is the classical Kolmogorov turbulence (e.g., San & Maulik, 2018). A turbulent eddy whose scale is the Larmor radius of the electrons, $r_{L}=\gamma m_{e}c^{2}/(eB_{\rm dis})$, scatters the electrons. We estimate the diffusion timescale as $t_{\rm esc}=R_{j}^{2}/D_{R}$, where $D_{R}\approx(4/\pi)r_{L}c(R_{j}/r_{L})^{2/3}(B/\delta B)^{2}$ is the diffusion coefficient (see also Kulsrud, 2005).
Appendix E Magnetization parameter for power-law tail distribution